How much heat (in kJ) is required to warm 175 g of sand from 25.0oC to 75.0oC
Answer #1

The equation to find the heat which is being transferred to the sand is

Q = m * Cp * (Tf - Ti)

Q is the heat,

m is the mass ,

Cp is the heat capacity

Tf : final temperature

Ti i: initial temperature

heat capacity for the sand is 0.84 J / g C (from literature)

we have 175 g of sand,

Q = 175 * 0.84 * (75- 25) = 7350 Joules

Divide by 1000 to get value in kJ which is 7.35 KJ

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION